you invested some money into an account with fixed percent interest compound annually after 3 years your investment was worth $3275.29 after 5 years $3680.12 what is the annual interest rate? and what was the original amount invested? How many years will it take to double?
Answer by stanbon(75887)

you invested some money into an account with fixed percent interest compound annually after 3 years your investment was worth $3275.29 after 5 years $3680.12
-------------------
Using A(t) = P(1+(r/n))^(nt)
------------------------------------
what is the annual interest rate?
3680.12 = 3275.29(1+r)^5
rate = 0.0236
-----------------------------
and what was the original amount invested?
P = 3275.29
---------------------------
How many years will it take to double?
Solve: (1.0236)^t = 2
t = ln(2)/ln(1.0236) = 30 years when rounded up
===================
Cheers,
Stan H.
--------